# Representational Analysis of Binding in Large Language Models

### v1

@@ Abstract
Entity tracking is essential for complex reasoning. To perform in-context entity tracking, language models (LMs) must bind an entity to its attribute (e.g., bind a container to its content) to recall attribute for a given entity. For example, given a context mentioning “The coffee is in Box Z, the stone is in Box M, the map is in Box H”, to infer “Box Z contains the coffee” later, LMs must bind “Box Z” to “coffee”. To explain the binding behaviour of LMs, Feng and Steinhardt (2023) introduce a Binding ID mechanism and state that LMs use a abstract concept called Binding ID (BI) to internally mark entity-attribute pairs. However, they have not directly captured the BI determinant information from entity activations. In this work, we provide a novel view of the Binding ID mechanism by localizing the prototype of BI information. Specifically, we discover that there exists a low-rank subspace in the hidden state (or activation) of LMs, that primarily encodes the order of entity and attribute and which is used as the prototype of BI to causally determine the binding. To identify this subspace, we choose principle component analysis as our first attempt and it is empirically proven to be effective. Moreover, we also discover that when editing representations along directions in the subspace, LMs tend to bind a given entity to other attributes accordingly. For example, by patching activations along the BI encoding direction we can make the LM to infer “Box Z contains the stone” and “Box Z contains the map”. $\ast$ Equal contribution
@@
Representational Analysis of Binding in Large Language Models
@@ 1 Introduction
The ability of a model to track and maintain information associated with an entity in a context is essential for complex reasoning Karttunen (1976); Heim (1983); Nieuwland and Van Berkum (2006); Barzilay and Lapata (2008); Kamp et al. (2010). To recall attribute information in the context, the model must bind entities to their attributesFeng and Steinhardt (2023). For example, given Sample 1 and 1, a model must bind the entities (e.g., “Box Z”, “Box M”, “Box H”, “Alex”, “John” and “Carl”) to their attributes (e.g., “coffee”, “stone”, “map”, “bean”, “pie” and “fruit”) respectively so as to accurately recall such as what is in “Box Z” or what is sold by “Alex” without confusion. Binding has also been studied as a fundamental problem in Psychology Treisman (1996). To uncover how Language Models (LMs) realize binding in term of internal representation, Feng and Steinhardt (2023) introduce a Binding ID mechanism and state that LMs apply a abstract concept called Binding ID (BI) to bind and mark Entity-Attribute (EA) pairs (e.g., “Box Z” and “coffee”, as shown in Sample 1, where BI is denoted as a numbered square). They also claim that the BI is represented as a vector to be added on the representation (or activation) of a EA pair so that the common vector is used as a key clue to search attribute for a given entity. However, they have not directly captured the prototype of BI information (called BI information hereafter for simplicity) from the activations.
@@
Figure 1: Our main finding on Binding ID (BI) subspace intervention. Patching entity (e.g., "Z") representations along BI direction in activation space yields corresponding changes in model output.
@@
Since binding is the foundational skill that underlies entity tracking Feng and Steinhardt (2023), in this work, we take the entity tracking task Kim and Schuster (2023); Prakash et al. (2024) as a benchmark to evaluate the LM’s binding behaviour. Based on the analysis of internal representation on this task, we localize the BI information from the activiations and provide a novel view of the Binding ID mechanism. Specifically, we discover that LMs encode (or store) BI information into a low-rank subspace (called BI subspace hereafter) that primarily encodes the order (i.e., the input order from left to right) of entity and attributes and which is used as the prototype of BI to determine binding. To identify the BI subspace, we take Principle Component Analysis (PCA) as our first attempt  [^1] Besids PCA, we also attempt partial least squares regression for capturing BI subspace. Since they achieve similar regression score, we adopt PCA for simplicity. See Appendix (§A.2) for details. to capture the subspace representing the BI information, and it is empirically proven be effective. Therefore, our findings confirm and extend prior understanding of BI in LMs.
Going beyond the prior work on BI mechanism Feng and Steinhardt (2023), we show that by causally intervening along the BI encoding Principle Component (PC), LMs swap the binding and infer a new attribute for a given entity accordingly. That is, we find a consistent causal relationship between the BI subspace intervention and the inferred attributes by LMs. For example, as shown in Figue 1, by patching activations along a direction (i.e., PC1), we can make the LMs to infer "Box Z contains the stone" and "Box Z contains the map" instead of "Box Z contains the coffee".
@@
Overall, our findings suggest that LMs encodes BI information into a subspace of entity activations that primarily encodes the appearance order (or reversed order) of a EA pair in a given context. Moreover, the discovered BI subspace plays a crucial role in the in-context binding computation. In addition, we find that BI subspace not only exists in the Pretrained large LMs such as Llama2 Touvron et al. (2023) and Llama3 AI@Meta (2024), but also in the code fine-tuned LM such Float-7B Prakash et al. (2024).
2 Finding Low-rank Subspace for Binding ID
@@
In this section we describe our Principle Component Analysis (PCA) based method to locate the BI subspace in activations of LMs. Firstly we extract entity activation from LMs as shown in Figure 1. Given a LM (e.g., Llama2), and a collection of texts which describe a set of EA pairs for a relation such as Sentence 1 for relation “is_in”, we extract the activation of entity token (e.g., “Z”) in query (denoted as $\mathbf{x}_{i}$) from certain layer  [^2] See Appendix (§A.3) for the layer selection. and construct a activation matrix $M_{r}\in R^{n\times d}$ for a relation $r$, where $n$ denotes the number of entities and $d$ denotes the dimension of the activation. The row $i$ of $M_{r}$ is the activation of an entity token (i.e., $\mathbf{x}_{i}$).
@@
PCA has been applied for identifying various subspace (or direction) such as the subspace encoding language bias Yang et al. (2021), truth value of assertions Marks and Tegmark (2023) and sentiment Tigges et al. (2023). Inspired by these studies, we choose PCA as our first attempt to localize BI subspace. We hypothesize that in a activation subspace, entities with the same BI tend to cluster together (w.r.t the ones with different BIs), even though these entities have different semantic meaning, and the BIs are encoded as directions (or a PC) in the subspace. For convenience, we number BIs in left-to-right order, and the leftmost BI$=0$.
@@
To capture the subspace, or BI direction, we leverage PCA, which identifies the principle directions of a space. Specifically, the PCA of a activation matrix is ${M_{r}}=U_{r}\Sigma_{r}V^{T}_{r}$, where the columns of $V_{r}\in R^{d\times d}$ are principle directions of $M_{r}$. We takes first $c$ columns of $V_{r}$ as the BI direction, denoted as $B_{r}\in R^{d\times c}$.
3 BI Subspace Visualization
@@
Figure 2: Layer-wise BI subspace visualization on Llama2-7B.
@@
We adopt a subset of the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024), which contains $n=1000$ samples, to create layer-wise activation matrix $M_{r}^{l}$. We then uses the $M_{r}^{l}$ to extract the layer-wise BI subspace projection matrix $B_{r}^{l}\in R^{d\times 2}$ to visualize the activation. Figure 2 shows the embedding visualization on Llama2-7B, where each point represents the activation of an entity projected via the $B_{r}^{l}$, and the colors represent BIs. From which, we can observe that middle layers, such as layer 8, have a clearly visible direction along which BI increases (or decreases), while the others have tangled distribution. We also observe similar property of distribution on Llama3-8B and Float-7B (§A.4). This indicates that LMs use the middle layers to encode BI information. We call this dimension that represents the order of BI as BI Principle Component (BI-PC). In the following section, we apply causal intervention on the BI-PC to analyze how BI-PC affect the model output.
4 Causal Interventions on Binding ID Subspace
@@
Table 1: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: is_in”, where color denotes the BI.
By projecting the activation matrix $M_{r}$ into the BI subspace, we have found a correlative evidence for the existence of the direction (i.e., BI-PC) that encodes Binding IDs. However, it is possible that the BI information is encoded in the BI subspace but has no effect on model output.
In order to test if Binding IDs are not only encoded in the BI subspace, but that these representations can be steered so as to swap the binding and change LM’s output. We now perform interventions to establish the causality. That is, we want to find out if making BI swapping interventions leads to a change in model output.
@@ 4.1 Activation Patching
Since LM computation graph could be viewed as causal graph Meng et al. (2022); McGrath et al. (2023), we intervene on model activations via activation patching Vig et al. (2020); Wang et al. (2022); Zhang and Nanda (2023); Heinzerling and Inui (2024); Engels et al. (2024), and observe the effect on model output. Unlike the common activation patching setup in which one replaces activations resulting from an original input with activations from a corrupted one, we create patches by editing activations along a particular direction (i.e., along BI-PC), similar to the activation editing method of Matsumoto et al. (2023); Heinzerling and Inui (2024); Engels et al. (2024). Although automatic methods for localizing model circuits of interest have been proposed Conmy et al. (2023); Kramár et al. (2024), for simplicity we perform a coarse layer-wise search based on the effect of activation patching in a development set, as shown in Appendix (§A.3), and use the found setting for all experiments.
@@ 4.2 Setting
Dataset To explore the internal representation that enables binding, we adopt the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024). The dataset contains English sentence describing a set of objects located in a set of boxes with difference labels, and the task is to infer what is inside a given box. For instance, when a LM is presented with “The coffee is in Box Z, the stone is in Box M, the map is in Box H, … Box Z contains the”, it should infer the next token as “coffee”. Each sample involves $7$ AE pairs. To evaluate the binding in various context, we also apply the templates shown in Table 2 to generate other $5$ datasets with different relation, where $a_{i}$ and $e_{i}$ denotes the attribute and entity, and they are sampled from a fixed pool of $224$ one-token objects (e.g., “dog”, “corn” and “cookie”) and $523$ of one-token names (e.g., “Alex”, “Juli” and “Dan”) respectively. We sample $n=1000$ context from each dataset to run the following analysis.
@@
Metrics We apply two evaluation metrics. The logit difference metric introduced in Wang et al. (2022), which calculates difference in logits between original and intervened answers, as well as the "logit flip" accuracy metric Geiger et al. (2022), which represents the proportion of cases where we alter the model output after a causal intervention.
@@
Figure 3: Logit Difference (LD) for BI-PC based intervention across datasets on Llama2-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the LD, BI_i represents each target attribute and the light yellow bottom line indicates the LD of original attribute (i.e., $a_{0}$). Here, $l=8$, $v=2.5$, and $\alpha=3.0$.
4.3 Results: Direct Editing BI Subspace
We hypothesize that LMs encode BI information into a low-rank BI subspace. Therefore, we wonder if a LM changes the binding behavior, when adding a particular value $v$ (called step hereafter) along the BI-PC mentioned in Section (§2). For example, if we add one unit of $v$ on the activation of $e_{0}$, the LM will reset its BI as $1$ and bind attribute $a_{1}$ to entity so that infers the $a_{1}$ as the attribute of $e_{0}$ instead of the original $a_{0}$. Similarly, adding two units of $v$ will make the LM infer $a_{2}$ for $e_{0}$, and so on. We intervene via the Equation 1, where $\mathbf{x}_{0,l}$ is the original activation of $e_{0}$ (i.e., the leftmost entity) in layer $l$, $\mathbf{x}_{0,l}^{*}$ is the intervened activation, $B_{r}$ is the BI subspace projection matrix mentioned in Section (§2), $\alpha$ is a hyper-parameter to scale the effect of intervention and $\beta$ ($0\leq\beta\leq 6$) denotes the number of steps.
@@ 4.3 Results: Direct Editing BI Subspace
Table 1 lists several examples under the BI subspace intervention on the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024). We also list the examples from other datasets in Appendix (§A.6). We can see that when adding $1$ step along BI-PC, the model selects “stone” for entity “Z” instead of its original attribute “coffee”. Similarly, when the step is doubled, the model will select attribute “map” for the entity, and so on. Although the attribute selection does not strictly follow the number of steps, this indicates, to some extent, that changing the value along BI-PC can induce the swap of attribute.
Besides the qualitative analysis, we also conduct quantitative analysis for the causality between the BI subspace based activation patching and the binding behaviour of the LM. We plot mean-aggregated effect of directed activation patching across multiple datasets in Figure 3. Figure 3 indicates how the Logit Difference (LD) of each attribute changes as the step increases. We can observe that as the number of steps increases, LD of the original attribute decreases. In contrast, LD of other attributes gradually increase until a certain point and then gradually decrease. Given a candidate attribute, its peak point generally corresponds the number of steps that equals to its BI. For instance, when adding $3$ steps, the points of BI_3 (i.e., attributes of BI$=3$) on step$=3$ achieve the highest LD score. This indicates that by adjusting the value along the BI-PC, we can increase the probability of the corresponding attribute, thereby swap the answer.
Similarly, Figure 4 illustrates the relation between the number of steps and the logit flip, which gauges the percentage of the predicted attributes under an intervention. Figure 4 shows that as the step increases, the proportion bar becomes darker, it means that the model promotes the proportion of the corresponding following attribute in its inference. For instance, when adding $3$ step on the subspace, the $a_{3}$ (i.e., BI_3) becomes the major of the answers. This proves that the discovered subspace stores BI information, and the subspace will causally affect the computation of Binding in a LM. See Appendix (§A.9) for the results on Llama3-8B.
@@
Figure 4: Logit flip for BI-PC based intervention across datasets on Llama2-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the proportion of each inferred attribute in model output.
4.4 Results: Activation Steering on BI Subspace

### v3

# Representational Analysis of Binding in Language Models

@@ Abstract
Entity tracking is essential for complex reasoning. To perform in-context entity tracking, language models (LMs) must bind an entity to its attribute (e.g., bind a container to its content) to recall attribute for a given entity. For example, given a context mentioning “The coffee is in Box Z, the stone is in Box M, the map is in Box H”, to infer “Box Z contains the coffee” later, LMs must bind “Box Z” to “coffee”. To explain the binding behaviour of LMs, Feng and Steinhardt (2023) introduce a Binding ID mechanism and state that LMs use a abstract concept called Binding ID (BI) to internally mark entity-attribute pairs. However, they have not captured the Ordering ID (OI) from entity activations that directly determines the binding behaviour. In this work, we provide a novel view of the BI mechanism by localizing OI and proving the causality between OI and binding behaviour. Specifically, by leveraging dimension reduction methods (e.g., PCA), we discover that there exists a low-rank subspace in the activations of LMs, that primarily encodes the order (i.e., OI) of entity and attribute. Moreover, we also discover the causal effect of OI on binding that when editing representations along the OI encoding direction, LMs tend to bind a given entity to other attributes accordingly. For example, by patching activations along the OI encoding direction we can make the LM to infer “Box Z contains the stone” and “Box Z contains the map”. The code and datasets used in this paper are available at https://github.com/cl-tohoku/OI-Subspace.
@@
Representational Analysis of Binding in Language Models
@@ 1 Introduction
The ability of a model to track and maintain information associated with an entity in a context is essential for complex reasoning Karttunen (1976); Heim (1983); Nieuwland and Van Berkum (2006); Barzilay and Lapata (2008); Kamp et al. (2010). To recall attribute information for a given entity in a context, the model must bind entities to their attributes Feng and Steinhardt (2023). For example, given Sample 1 and 1, a model must bind the entities (e.g., “Box Z”, “Box M”, “Box H”, “Alex”, “John” and “Carl”) to their corresponding attributes (e.g., “coffee”, “stone”, “map”, “bean”, “pie” and “fruit”) so as to recall (or answer) such as what is in “Box Z” or what is sold by “Alex” without confusion. Binding has also been studied as a fundamental problem in Psychology Treisman (1996).
Figure 1: Our main finding on Ordering ID (OI) subspace intervention. Patching entity (e.g., "Z") representations along OI direction (i.e., PC1) in activation space yields corresponding changes in model output.
To uncover how Language Models (LMs) realize binding in term of internal representation, Feng and Steinhardt (2023) introduce a Binding ID mechanism and state that LMs apply a abstract concept called Binding ID (BI) to bind and mark Entity-Attribute (EA) pairs (e.g., “Box Z” and “coffee” in Sample 1, where BI is denoted as a numbered square). They also claim that the BI is represented as a vector to be added on the representation (or activation) of a EA pair so that the common vector is used as a key clue to search attribute for a given entity. However, they have not captured the Ordering ID (OI) information from the entity (or attribute) activations that causally affects binding behaviour and thus BI information as well. Here, OI is defined as the input order (or ordering index) of entities and attributes, no matter they are bound by a relation (e.g., “is_in” in Sample 1) or not, such as the indexing number in Sample 1 and Sample 1. We can observe that in a 1E-to-1A bound context, such as in Sample 1, BI and OI are interchangeable.
@@
Since binding is the foundational skill that underlies entity tracking Feng and Steinhardt (2023), in this work, we take the entity tracking task Kim and Schuster (2023); Prakash et al. (2024) as a benchmark to analyze the LM’s binding behaviour. Based on the analysis of internal representation on this task, we localize the OI information from the activiations and provide a novel view of the BI mechanism. Specifically, we apply Principle Component Analysis (PCA) as well as other dimension reduction methods such as Independent Component Analysis (ICA)  [^1] See Appendix (§A.2) and Appendix (§A.3) for details. to analyze the activations of LMs, and which are empirically proven to be effective. We discover that LMs encode (or store) the OI information into a low-rank subspace (called OI subspace hereafter), and the discovered OI subspace can causally affect binding behaviour and thus BI information as well. That is, we find that by causally intervening along the OI encoding Principle Component (PC), LMs swap the binding and infer a new attribute for a given entity accordingly. For example, as shown in Figure 1, by patching activations along the direction (i.e., PC1), we can make the LMs to infer “Box Z contains the stone” and “Box Z contains the map” instead of “Box Z contains the coffee”. Therefore, our findings extend the previous BI based understanding of binding in LMs Feng and Steinhardt (2023) by revealing the causality between OI and binding.
@@
Overall, our findings suggest that LMs encode OI information into a subspace of LMs’ activations that primarily encodes the order index of entities and attributes in a given context. What is more, the discovered OI subspace plays a crucial role in the in-context binding computation. In addition, we find that such OI subspace that determines binding is prevalent across multiple LM families such as Llama2 Touvron et al. (2023) (and Llama3 AI@Meta (2024)), Qwen1.5 Bai et al. (2023) and Pythia Biderman et al. (2023), and the code fine-tuned LM Float-7B Prakash et al. (2024).
2 Finding OI Subspace
@@
In this section we describe our Principle Component Analysis (PCA) based method to localize the OI subspace in activations of LMs. As shown in Figure 1, we firstly extract entity activation from LMs. Given a LM (e.g., Llama2), and a collection of texts which describes a set of EA pairs related by a relation such as “is_in” in Sample 1, we extract the activation of entity token (e.g., “Z”) in query (denoted as $\mathbf{x}_{i}$) from certain layer  [^2] See Appendix (§A.4) for the layer selection. and construct a activation matrix $M_{r}\in R^{n\times d}$ for a relation $r$, where $n$ denotes the number of entities and $d$ denotes the dimension of the activation. The row $i$ of $M_{r}$ is the activation of an entity token (i.e., $\mathbf{x}_{i}$).
PCA has been applied for identifying various subspace (or direction) such as the subspace encoding language bias Yang et al. (2021), truth value of assertions Marks and Tegmark (2023) and sentiment Tigges et al. (2023). Inspired by these studies, we choose PCA as our first attempt to localize OI subspace. In addition, we also apply other dimension reduction methods such as Independent Component Analysis (ICA) to capture OI subspace, and which are empirically proven to be effective. See Appendix (§A.2) and Appendix (§A.3) for details.
@@
We hypothesize that in a activation subspace, entities with the same OI tend to cluster together (w.r.t the ones with different OIs), even though these entities have different semantic meaning, and the OIs are encoded as directions (or a PC) in the subspace. For convenience, we number OIs in left-to-right order, and the leftmost OI$=0$.
@@
Since PCA is applied to identify the principle directions from a multidimensional space, we leverage PCA to capture OI subspace (or direction) from activations of LMs. Specifically, the PCA of a activation matrix is ${M_{r}}=U_{r}\Sigma_{r}V^{T}_{r}$, where the columns of $V_{r}\in R^{d\times d}$ are principle directions of $M_{r}$. We takes first $c$ columns of $V_{r}$ as the OI direction, denoted as $B_{r}\in R^{d\times c}$.
3 OI Subspace Visualization
@@
Figure 2: Layer-wise OI subspace visualization on Llama2-7B, where “BI” primarily denotes OI.
@@
We adopt a subset of the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024), which contains $n=1000$ samples, to create layer ($l$) wise activation matrix $M_{r}^{l}$. We then uses the $M_{r}^{l}$ to extract the layer-wise OI subspace projection matrix $B_{r}^{l}\in R^{d\times 2}$ to visualize the activations. Figure 2 shows the embedding visualization on Llama2-7B, where each point represents the activation of an entity projected via the $B_{r}^{l}$, and the colors represent OIs. From which, we can observe that middle layers, such as layer 8, have a clearly visible direction along which OI increases, while the others have tangled distribution.
We also observe similar pattern of distribution on Llama3-8B, Float-7B (§A.5) and other LM families such as Qwen1.5 and Pythia (§A.6). This indicates that LMs use the middle layers to encode OI information, and the finding is prevalent across multiple LM families. This finding is also consistent with the “stages of inference hypothesis” Lad et al. (2024) stating that the function of early layers is to perform detokenization, middle layers do feature engineering, and late layers map the representations from the middle layers into the output embedding space for next-token prediction. According to the hypothesis, we would expect to find the ordering feature most prominently represented in middle layers, which is exactly what the visualization shows. We call this dimension that represents OI as OI Principle Component (OI-PC). In the following section, we apply causal intervention on the OI-PC to analyze how OI-PC affect the model output.
4 Causal Interventions on OI Subspace
@@
Table 1: Attributes inferred by Llama2-7B as a result of directed activation patching along OI-PC in the OI subspace on the dataset of “r: is_in”, where color denotes the BI.
By projecting the activation matrix $M_{r}$ into the OI subspace, we have found a correlative evidence for the existence of the direction (i.e., OI-PC) that encodes OI information. However, it is possible that the OI information is encoded in the OI subspace but has no effect on LMs’ binding behaviour.
In order to test if OIs are not only encoded in the OI subspace, but that these representations can be steered so as to swap the binding and change LM’s output, in this section, we perform interventions to analyze the causality. That is, we want to find out if making interventions along OI-PC leads to a change in LM’s binding computation.
@@ 4.1 Activation Patching
Activation Patching (AP) Vig et al. (2020); Geiger et al. (2020, 2021); Wang et al. (2022); Stolfo et al. (2023); Heinzerling and Inui (2024); Engels et al. (2024); Hanna et al. (2024) has been recently proposed to causally intervene computational graph of a LM so as to interpret the function of a target computational node (or edge). The process of AP usually involves preparing corrupted input, using its activations to replace the corresponding ones obtained from original input and analyzing the effect on model output. Different with the common AP setup, we realize AP by directly editing activations along a particular direction (i.e., along OI-PC), similar to the activation editing method of Matsumoto et al. (2023); Heinzerling and Inui (2024); Engels et al. (2024).
@@ 4.2 Setting
Dataset To explore the internal representation that enables binding, we adopt the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024). The dataset consists of English sentence describing a set of objects (here called attributes) located in a set of boxes with difference labels (here called entities), and the task is to infer what is contained by a given box. For instance, when a LM is presented with “The coffee is in Box Z, the stone is in Box M, the map is in Box H, … Box Z contains the”, the LM should infer the next token as “coffee”. Each sample involves $7$ EA pairs. To evaluate the binding in various context, we also apply the templates shown in Table 2 to generate other $5$ datasets with different relation, where $a_{i}$ and $e_{i}$ denotes the attribute and entity, and they are sampled from a fixed pool of $224$ one-token objects (e.g., “dog”, “corn” and “cookie”) and $523$ of one-token names (e.g., “Alex”, “Juli” and “Dan”) respectively. We sample $n=1000$ context from each dataset to run the following analysis.
@@
Metrics We apply two evaluation metrics: logit difference and logit flip. The logit difference metric is introduced in Wang et al. (2022), which calculates difference in logits of a target token between original and intervened setting. The "logit flip" accuracy metric is introduced by Geiger et al. (2022), which represents the proportion of candidate tokens in model output after a causal intervention.
@@
Figure 3: Logit Difference (LD) for OI-PC based intervention across datasets on Llama2-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the LD, BI_i represents each target attribute and the light yellow bottom line indicates the LD of original attribute (i.e., $a_{0}$). Here, $l=8$, $v=2.5$, and $\alpha=3.0$.
4.3 Results: Direct Editing OI Subspace
We hypothesize that LMs encode OI information into a low-rank subspace that causally affect BI information and thus binding behaviour as well. Therefore, we wonder if a LM changes the binding behavior, when adding a particular value $v$ (called step hereafter) along the OI-PC mentioned in Section (§2333where PC2 is assigned a fixed value, which is a hyperparameter.. For example, if we add one unit of $v$ on the OI-PC of $e_{0}$, the LM will reset its BI as $1$ and bind attribute $a_{1}$ to the entity based on the similarity of BI so that infers the $a_{1}$ as the attribute of $e_{0}$ instead of the original $a_{0}$. Similarly, adding two units of $v$ will make the LM infer $a_{2}$ for $e_{0}$, and so on. We intervene via the Equation 1, where $\mathbf{x}_{0,l}$ is the original activation of $e_{0}$ (i.e., the leftmost entity) in layer $l$, $\mathbf{x}_{0,l}^{*}$ is the intervened activation, $B_{r}$ is the OI subspace projection matrix mentioned in Section (§2), $\alpha$ is a hyper-parameter to scale the effect of intervention and $\beta$ ($0\leq\beta\leq 6$) denotes the number of steps.
@@ 4.3 Results: Direct Editing OI Subspace
Table 1 lists several examples under the OI subspace intervention on the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024). We also list the examples from other datasets in Appendix (§A.7). We can see that when adding $1$ step along OI-PC, the model selects “stone” for entity “Z” instead of its original attribute “coffee”. Similarly, when the step is doubled, the model will select attribute “map” for the entity, and so on. This indicates that changing the value along OI-PC can induce the swap of attribute. In addition, we notice that some attributes are repeated or skipped after the intervention. The reason is that, as shown in Figure 2, OI is represented as a continuous range (e.g., $[a_{0},b_{0}],[a_{1},b_{1}],[a_{2},b_{2}],...$) on OI-PC, implying the points in the same range share OI, and if a sample (e.g., its OI-PC is $s_{i}$) is still in the range (e.g., $a_{0}<(s_{i}+v)<b_{0}$) or skip its neighbouring range (e.g., $a_{2}<(s_{i}+v)<b_{2}$) after the intervention (e.g., $s_{i}+v$), its OI will be the same or added by $2$, and thus the binding information will change accordingly.
Besides the qualitative analysis, we also conduct quantitative analysis for the causality between the OI subspace based AP and the binding behaviour of LMs. We plot mean-aggregated effect of the OI-PC based AP across multiple datasets in Figure 3. Figure 3 indicates how the Logit Difference (LD) of each attribute changes as the step increases. We can observe that as the number of steps increases, LD of the original attribute decreases. In contrast, LD of other attributes gradually increase until a certain point and then gradually decrease. Given a candidate attribute, its LD peak roughly corresponds to the number of steps that is equal to its BI. For instance, when adding $3$ steps, the points of BI_3 (i.e., attributes of BI$=3$) achieve the highest LD score. This indicates that by adjusting the value along the OI-PC, we can adjust BI information and thus increase the logit score of the corresponding attribute.
Similarly, Figure 4 illustrates the relation between the number of steps and the logit flip, which gauges the percentage of the predicted attributes under an intervention. Figure 4 shows that as the step increases, the proportion bar becomes darker, it means that the model promotes the proportion of the corresponding attribute in its inference. For instance, when adding $3$ step on the subspace, the $a_{3}$ (i.e., BI_3) becomes the major of the answers. This proves that the OI-PC based interventions can causally affect BI information as well as the computation of Binding in a LM. See Appendix (§A.10) for the results on Llama3-8B, Appendix (§A.6) for the results on Qwen1.5-7B and Pythia-6.9B, and Appendix (§A.9) for the results on Float-7B.
@@
Figure 4: Logit flip for OI-PC based intervention across datasets on Llama2-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the proportion of each inferred attribute in model output.
4.4 Results: Activation Steering on OI Subspace